OpenStudy (anonymous):

mom bought 3 1/3 pounds of berry. she estimates she will need 1/9 pound of berries for each smoothie. how many smoothies can she make with the berries she bought? Thanks!!

OpenStudy (anonymous):

3 1/3 = 10/3 10/3 divded by 1/9 = 30 smoothies

OpenStudy (anonymous):

thanks